MEASUREMENT OF OH PRESSURE BROADENING

BRIAN J. DROUIN AND HERBERT M. PICKETT, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

Measurement of atmospheric signatures of the hydroxyl radical using radiometric techniques require knowledge of the temperature dependent pressure broadening in air. Utilizing the Balloon OH instrument; a side-band spectrometer that operates in the same band-pass as the THz microwave limb sounder on the Aura spacecraft has been constructed. Embedded in an all-vacuum apparatus including a 2.2 meter absorption cell and filtering optics the system allows spectroscopic measurements in the 2.5-2.54 THz, 225-300 K range. New measurements of the hydroxyl radical lineshapes at 2.509 and 2.514 THz are presented.